OpenStudy (anonymous):

a-(3a-6/6)=(3a+2/3) a=?

OpenStudy (anonymous):

a=1/15

OpenStudy (anonymous):

\[a-\left(\begin{matrix}3a-6 \\ 6\end{matrix}\right)=\left(\begin{matrix}3a+2 \\ 3\end{matrix}\right)\]